5. Period 1: Firm A (Pre-merger) (20 points) Assumptions: • Price = $4 per unit of output sold • Variable costs = $2.75 per u can anyone help me with Q d and e please? include work
Answer #1

Calculation as per the given data:

Pre-Merger

Price per unit

$4.00

Variable Costs per unit

$2.75

Fixed Costs

$1,000,000

No. of units per year

1,000,000

Plant capacity utilized

50.00%

Plant capacity

2,000,000

Post-Merger

Price per unit

$4.00

Variable Costs per unit

$2.75

Fixed Costs

$1,000,000

No. of units per year

1,500,000

Pre-Merger

Post-Merger

A

Sale Price per unit

$4.00

$4.00

B

Variable Cost per unit

$2.75

$2.75

C = A - B

Contribution per unit

$1.25

$1.25

D

No. of units

1,000,000

1,500,000

E = A x D

Total Sales

$4,000,000

$6,000,000

F = B x D

Total variable cost

$2,750,000

$4,125,000

G = E - F

Total contribution

$1,250,000

$1,875,000

H

Fixed Costs

$1,000,000

$1,000,000

I = G - H

Profit

$250,000

$875,000

J = I / E

Profit Margin

6.25%

14.58%

  1. The profit margin per dollar of revenue is that part of $1.00 which is earned by the company as profit. As calculated, the pre-merger margin is 6.25%. Hence, the profit margin per dollar is ($1.00 x 6.25%) i.e. $0.0625.

  1. As seen in the above table, the post-merger profit is $875,000.

  1. As seen in the above table, the post-merger profit margin is 14.58%.

  1. As calculated, the post-merger margin is 6.25%. Hence, the profit margin per dollar is ($1.00 x 14.58%) i.e. $0.1458.